Crew 211 Science Report 25Apr20189

[title Science Report – April 25th]

[category science-report]

Science Report 25 April 2019
Crew 211 – UCL to Mars
Submitted by Science Officer Maxime Bernard

1) Commander Carl Dahlqvist

Working on the software to allow the main module to receive data from the secondary modules.

2) XO Officer

Coding a trilateration algorithm to compute the position of a target from the respective distances from that target to the 3 antennas.

3) HSO Officer Benjamin Flasse

Every morning and every evening, the HSO still gather data about the mensurations, weight, brain awakeness, tension and pulse of the crew.

4) Crew Astronomer Eleonore Lieffrig

Working on the images she got from the remote telescope.
In the meantime she started to work on the Brownian motion operator.

5) Crew Engineer Julien Amalaberque

No major milestone reach for now.

6) Crew Biologist Chloé Peduzzi

She checked the spirulina through the microscope and confirmed it was the right species (Arthrospira Platensis). Moreover, she observed their spiral shape which indicate their healthy state.
She then proceeded to the aeration of all the spirulina culture.

7) Crew Journalist Nathan Pechon

We made a meeting during which he presented us law articles about birth and immigration rights on a martian colony and we had to vote these articles.

8) Crew Scientist Maxime Bernard

He assembled his detector and started the acquisition of data from muon interactions

Sol Summary – April 25th

Good evening Atila,

Please find below the Sol 4 Summary.

Crew 211 Sol 4 Summary Report 25-APR-2019

Sol: 4

Summary Title: Rescuing a Martian

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The day started with an EVA to explore the Special Region. Unfortunately the battery level of the Spirit rover decreased to fast so we couldn’t reach the Special region. After a short briefing with the Hab, we decided to explore instead the Galileo road 1104 for about 30 minutes. We then worked on our experiments until the lunch. After a nice lunch altogether, the HSO performed a quick check-up on the crew member who twisted his ankle three days ago. As the ankle was still pretty much swelled, we decided with Dr Rupert that it would be better to see a doctor. We therefore drove the crew car to the Bicknell Health Care center. Fortunately, the ankle was not broken and the crew member should be healed within a week. The rest of the team worked during the afternoon on their respective experiments.

Look Ahead Plan: An EVA is planned in the morning to explore Special Region. Half of the crew will stay at the base to work, as usual, on their experiments. The rest of the day will be devoted to these experiments.

Anomalies in work: No issues.

Weather: Hot, bright and sunny most of the day, but windy in the evening.

Crew Physical Status: Twisted ankle still bloated but should be healed within a week

EVA: The aim of today’s EVA was to explore the Special Region. Unfortunately the battery level of the Spirit rover decreased to fast so we couldn’t reach the Special region. After a short briefing with the Hab, we decided to explore instead the Galileo road 1104 for about 30 minutes.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Science Report, Operations Report, EVA report, EVA Request, Daily Photo Report

Support Requested: None

Crew 211 EVA Report 25Apr2019

[title EVA Report – April 25th]

[category eva-report]

Crew 211 EVA Report 25-04-2019

EVA #4

Author: Simon Collignon

Purpose of EVA: Explore the Special Region in order to explore and search for ancient martian life.
Start time: 09:00 am
End time: 11:00 am

Narrative: The aim of this EVA was to explore the Special Region as said in the purpose section of this EVA report. Unfortunately, the rover SPIRIT was discharging too fast, the battery’s autonomy was not high enough to reach the Special Region. Therefore, we decided to stop on Cow Dung Road at the intersection with Galileo Road 1104. After a short exchange with the Hab, we decided to take Galileo road and explore the area delimited by Cow Dung Road, Galileo Road, Cactus Road and Stream Bed Connector. Sadly, there were no sign of ancient Martian on this area. Next time we’ll try to reach Special Region

Destination: Galileo Road 1104

Coordinates (use UTM NAD27 CONUS): Latitude: 4252000, longitude: 519500 (Intersection of Stream Bed Connector and Cow Dung Road)

Participants: Carl-Henrik Dahlqvist (commander of the EVA), Simon Collignon (medical officer), Chloe Peduzzi, Nathan Pechon.

Road(s) and routes per MDRS Map: On Cow Dung Road from Hab to the intersection with Galileo Road 1104. Then we took Galileo road1104 and stopped 500 meters after.

Mode of travel: Driving and hiking

Greenhab Report – April 25th

Crew 211 GreenHab Report – 25-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 3 between 10:00 and 18:00 + fan turned up to 2 the rest of the time

Shade cloth: 80%

Average temperatures:

Low: 15°C

High: 46°C

Hours of supplemental light: 0

Daily water usage for crops: 3 gallons

Water in Blue Tank: 100 gallons

Time(s) of watering for crops: 11:30

Changes to crops: All plants look healthy. More flowers are visible on the lettuce (old seed). Sugar Snap peas, spinaches (planted 04/06) and sage grow very well.

Narrative: Still hot today despite the fan was turned up to 3. For our dinner we harvested parsley and basil (from the GreenHab).

Harvest: 13g basil + 9g parsley

Support/supplies needed: None

Journalist Report – April 25th

Good evening Atila,

Today was our fourth simulation day. We woke up at 7:30 am and we had breakfast at 8:00. Yesterday, Maxime and Eléonore made bread, we ate it, it was delicious and not dry. We spread peanut butter and Carl’s jam on it.

Carl, Simon, Chloé and Nathan went to EVA at 9 am. They had to go to Special Region to see dinosaurs bones but they had a problem with the battery of one of the rovers. Finally, they went to Cactus Road.

In the same time, the others cleaned the Hab and wrote a Facebook publication.

At 12 am, Benjamin and Nathan prepared the lunch. We ate pasta with tomatoes and mozzarella. It was cheesy. After meal, we discussed about Nathan’s experience and his martian constitution. We spoke about the political regime to introduce and reproduction on Mars.

At 2pm, the Commander, the HSO and the injured person went to clinic. Fortunately, he will get better soon. Maxime baked bread today. The rest of the crew worked on his experiences. Chloé observed spiruline with microscope.

At 6 pm, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. After the Cap Com, we will eat a soup with bread.

Like usual, we will do a crew meeting to speak about the day.

Operations Report – April 24th

Crew 211 Operations Report 24-APR-2019

SOL: 3

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10:30

On (last night): 8:30 pm

Off (this morning): 7 am

Solar— SOC 100 % at 20:10 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~240 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 105 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 6 units

Toilet tank emptied: yes

Deimos rover used: no (rover is out of order according to Director)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 89.0h

Beginning charge: 100

Ending charge: 84.4

Currently charging: yes

Opportunity rover used: yes

Hours: 57.6

Beginning charge: 95

Ending charge: 74

Currently charging: no

Curiosity rover used: yes

Hours: 90.8h

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Notes on rovers: Yesterday during the EVA the rover Opportunity suffered overheating and could not be started back. This morning’s mission was to bring it back so that the mechanics could fix it.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Some basil and tomatoes were harvested for cooking.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Simon finished his UWB stands and Carl continued assembling more probes during the afternoon after children finished visiting us.

Summary of any observatory issues: The Astronomer’s computer encounters sometimes freeze and displays artifacts on the screen. This might be the beginning of a GPU failure.

Summary of health and safety issues: The member with the twisted ankle feels better and can walk, but for extra safety HSO does not recommend EVA.

Questions, concerns and requests to Mission Support:

* We had an Internet problem last night which prevented us to send the operations report to CapComm. It solved itself at around 10am this morning.

* We overestimated the static water tank capacity, and after checking the level properly with a rope and measuring volumes we reestimated its current level. We will try to diminish consumption as much as we can, but it is already very low and we are investigating other reasons.

* The water pump leak is at a plastic joint. To fix it we would need to stop the waterflow and replace the joint with a new one. We will try to document further the following days. Attaching a picture of the pump with the mail (faulty joint is in red, and you can see that water has been leaking for it).

Copyright © The Mars Society. All rights reserved. | Main Site